# SIToBI - A Speech Prosody Annotation Tool for Indian Languages

### v1

@@ Abstract
The availability of prosodic information from speech signals is useful in a wide range of applications. However, deriving this information from speech signals can be a laborious task involving manual intervention. Therefore, the current work focuses on developing a tool that can provide prosodic annotations corresponding to a given speech signal, particularly for Indian languages. The proposed Segmentation with Intensity, Tones and Break Indices (SIToBI) tool provides time-aligned phoneme, syllable, and word transcriptions, syllable-level pitch contour annotations, break indices, and syllable-level relative intensity indices. The tool focuses more on syllable-level annotations since Indian languages are syllable-timed. Indians, regardless of the language they speak, may exhibit influences from other languages. As a result, other languages spoken in India may also exhibit syllable-timed characteristics. The accuracy of the annotations derived from the tool is analyzed by comparing them against manual annotations and the tool is observed to perform well. While the current work focuses on three languages, namely, Tamil, Hindi, and Indian English, the tool can easily be extended to other Indian languages and possibly other syllable-timed languages as well.
Index Terms:
Prosody Annotation, Tones and Break Indices, Speech Prosody, Segmentation
I Introduction
@@
Speech prosody plays a vital role in several applications and can be instrumental in improving human-computer interactions.While recent advancements in artificial intelligence have resulted in high-performing speech-based systems, these still do not accommodate prosodic variations in spotaneous speech or discourse. For instance, current text-to-speech (TTS) synthesizers produce speech that is highly intelligible and natural. However, these require a huge amount of training data and are still devoid of emotion and adequate prosodic variations. Incorporating prosodic information while training a TTS system could possibly improve the naturalness of synthetic speech, without necessitating an increase in the amount of training data. Another application that can benefit from the availability of prosodic information is speech-to-speech (S2S) translation. An S2S system involves an automatic speech recognition (ASR) system that provides text corresponding to a given speech signal. This text is then translated to the desired language, which is then synthesized using a TTS system. Incorporation of prosodic information along with the ASR output could aid in producing speech in the target language that carries the same sentiment as the speech in the source language. This is precisely where we have a problem. Also when the word order changes, incorporation of prosody in the target language is a big challenge. Further, each language has a unique prosodic structure and hence prosodic features could aid in language identification and in the incorporation of language specific prosody in TTS and S2S systems as well.
@@
While prosody is immensely useful in a wide range of applications, obtaining accurate prosodic annotations is usually a laborious task requiring manual intervention. Further, while a standard representation of prosody exists for English, namely ToBI (Tones and Break Indices) [1, 2], as well as for certain other languages, no such standard has been established for Indian languages. Therefore, the current work extends on the existing ToBI standard while incorporating additional information, namely time-aligned phonetic, syllabic, and word transcriptions and a relative intensity index, to develop an SIToBI (Segmentation with Intensity, Tones and Break Indices) tool for prosodic annotation.
@@
The ToBI system is a set of conventions for transcribing and annotating the prosody of speech, designed originally for English. Since the manual annotation of prosody is laborious, certain automatic labeling tools have been developed for English. One such tool is the Automatic ToBI (AuToBI) [3], which uses machine learning to predict ToBI labels from acoustic features of the speech signal. Another commonly used platform for annotation is praat [21]. Although originally designed for manual annotation, it hosts several plugins and scripts supporting semi-automatic prosodic annotation [4]. Web MAUS [5] is a web-based service for automatic segmentation and labeling of speech, including prosodic features. Another popular tool is PyToBI [6], which provides a Python interface for automatic labeling with ToBI. However, these are specifically designed for English.
@@
Efforts to study prosody in Indian languages—characterized by their syllable-centric structures—have largely focused on analyzing intonation, stress, and rhythm in languages such as Hindi, Bengali, Tamil, Kannada, and Telugu [12, 13, 14]. However, a standardized framework for representing prosodic features across these languages is yet to be established. Recognizing the similarities in their prosodic structures, this work proposes a common prosodic annotation standard and tool. Extending the ToBI (Tones and Break Indices) standard, the SIToBI tool automates the annotation of pitch contours and break indices for Indian languages, while incorporating additional features, including:
@@
The inclusion of these additional features enhances the utility and scope of the tool. Phoneme, syllable, and word boundaries enable precise alignment of prosodic features such as pitch and stress with linguistic units. They also significantly enhance speech synthesis and recognition systems, improving the naturalness and accuracy of text-to-speech (TTS) systems and automatic speech recognition (ASR) models.
The relative intensity index, on the other hand, offers crucial insights into stress and emphasis, contributing to more detailed prosodic modeling. This feature is especially useful for analyzing speaker emotion, intent, and conversational styles, enriching applications in areas such as emotions classification and expressive speech synthesis.
By combining tones and break indices with these additional features, SIToBI addresses limitations in existing prosody annotation tools, offering a more comprehensive framework for prosodic analysis. While designed for adaptability across Indian languages, the current implementation focuses on Tamil, Hindi, and Indian English, laying a robust foundation for broader applications in the future.
The paper is organized as follows: Section II provides a detailed description of the proposed tool, Section III evaluates the accuracy of the annotations provided by the tool, Section IV presents the use of the tool to perform language identification, and Section V concludes the paper.
II The SIToBI Tool
The SIToBI tool is designed to provide rich prosodic information from a speech signal. Given a speech signal, the tool converts it to a single channel and resamples it to a sampling frequency of 16 kHz. It then uses an ASR to obtain the corresponding orthographic transcription. This transcription, along with the speech signal, is used to derive syllable and word-level boundaries. The tool also processes the speech signal to estimate the relative intensity index, pitch contour labels, and break indices. The extracted information is then plotted as shown in Fig 1.
Figure 1: Annotations from the SIToBI Tool for the English sentence, ”For the twentieth time that evening the two men shook hands.”
In the SIToBI tool, an automatic speech recognition (ASR) system is used when only the speech signal (WAV format) is available, without accompanying text. In this scenario, the ASR generates an orthographic transcription from the audio. The other parameters are extracted as mentioned above. However, it’s important to note that the transcription produced by the ASR may sometimes be inaccurate, potentially leading to errors in the segmentation of the speech signal and in the extraction of other features.
@@ II The SIToBI Tool
Importantly, in both cases—whether the ASR system is used (when only audio is available) or not (when both audio and text are provided)—the SIToBI tool consistently applies language-independent models.
Considering the number of languages in India, building systems for each language individually is a complex and time-consuming task. For instance, creating a Text-to-Speech (TTS) system for a new language requires a thorough understanding of its phonotactics (sound structure), letter-to-sound (LTS) rules, and other linguistic elements. This process involves considerable effort and often requires the expertise of linguists familiar with the language. As Prakash et al. (2014) highlighted in their work, ”An approach to building language-independent text-to-speech synthesis for Indian languages” [19],prakash2023tts, building such systems from scratch for each language involves challenges such as training Hidden Markov Models, developing speech context modeling, and collecting language-specific data. These tasks are not only complex but also time-consuming.
By adopting language-independent models, the SIToBI tool eliminates the need for such language-specific efforts. These models offer the flexibility of adapting to new languages with minimal modification, making the process more efficient. They do not require the extensive linguistic resources that traditional, language-dependent models need. Instead, they rely on generalizable algorithms that can be easily applied across multiple languages, greatly improving the scalability and applicability of the SIToBI tool for the multilingual context of India.
@@
The block diagram 2 outlines the overall process of text and speech processing, where input speech and text are segmented and are processed via prosody modules. This sequential flow enables extraction of phoneme, word, and syllable segments, along with prosodic features, resulting in a comprehensive annotation of the input speech.
Figure 2: Block diagram
II-A Automatic Speech Recognition
To obtain the orthographic transcription corresponding to a given speech signal, the speech signal and the language (obtained from the user) are given to an ASR. In the current work, a pre-trained Data2Vec-aqc-based ASR model [15], trained on 30,000 hours of speech data from 24 Indian languages, including Tamil, Hindi, and Indian English, is employed.
II-B Segmentation
Once the text corresponding to a given speech signal is obtained from the ASR, it is used to generate time-aligned phonetic, syllabic, and word transcriptions. To achieve this, context-independent hidden Markov models (HMMs) are trained for each phoneme across the three languages considered  [^1] Due to the requirement for less data and faster training, we favor HMM-based training over neural network methods.. These HMMs are trained using three states, with one to five mixture components per state, depending on the number of available examples for each phoneme. The forced-Viterbi alignment procedure [16, 20] is initially applied to identify phoneme boundaries, and this process is repeated over four iterations to obtain the final boundaries.This iterative process ensures the accurate segmentation of speech data, enabling robust phoneme, syllable, and word-level boundary derivation. Monophone HMMs for all phonemes are generated using the label files obtained. With these models, forced-Viterbi alignment is performed iteratively to segment the rest of the data, as described below:
@@ II-B Segmentation
This iterative process ensures the accurate segmentation of speech data, enabling robust phoneme, syllable, and word-level boundary derivation.Since the tool is designed to support all three languages, language-independent phoneme models are also trained, where acoustically similar phonemes across the languages share the same model.
Syllabification involves combining phoneme boundaries to define syllable boundaries using a set of predefined rules. In this process, each vowel acts as the nucleus of a syllable, with consonants appearing at the beginning or end. The phoneme sequence is grouped into syllables in forms such as $V$, $C^{*}V$, $VC^{*}$, and $C^{*}VC^{*}$, where consecutive consonants are resolved based on language-specific phonological patterns. This systematic approach ensures precise syllable segmentation, facilitating subsequent linguistic and prosodic analysis.
While deriving the phoneme boundaries, it is ensured that there are silences inserted at the end of each word, though their duration might be close to zero at several places. These silences are used to combine the phoneme/syllable boundaries to derive word boundaries as discussed in section II
II-C Computation of the Relative Intensity Index
The tool then calculates the relative intensity index at the syllable level. The speech signal is divided into overlapping frames of 20 ms with a 10 ms hop length, and the Short-term energy (STE) is computed for each frame. The energy is subsequently normalized across the entire utterance. Based on the normalized energy, a relative intensity index with difference of 0.2 is assigned to each syllable on a scale from 1 to 5 as follows:
@@ II-C Computation of the Relative Intensity Index
Fig. 1(b) depicts the $RII$, where the syllable, “mehn” has the highest RII of 5, while the syllables, “dha”, “iyv”, “ningh”, “shuck”, “haendz”, etc. have the lowest RII of 1.
II-D Estimation of Break Indices
In order to determine the break indices, speech vs. silence discrimination is first performed using spectral flatness, which is a measure of the energy distribution across frequencies. Spectral flatness (SF) is computed frame-wise using the following equation, with the same frame length and hop length mentioned in the previous section.
@@ II-D Estimation of Break Indices
Break $\text{Break Index}=\begin{cases}1,&\text{if }l<80\text{ ms}\\ 2,&\text{if }80\text{ ms}\leq l<290\text{ ms }\\ 3,&\text{if }l\geq 290\text{ ms }\end{cases}$ (3)
II-E Labeling Pitch Contours
@@
Once the pitch contour, i.e., the time-varying $F_{0}$, is estimated, it undergoes syllable-level smoothing to reduce local fluctuations and abrupt changes. This is achieved using a polynomial of order 3, which fits a smooth curve to the pitch values over time, yielding a more natural representation of prosodic variations. Initially, eleven shapes of the pitch contour, namely L, H, HLL, HHL, LLH, LHH, HLH, LHL, hat, bucket, and flat, are identified, where L (low) represents a falling contour and H (high) represents a rising contour. To capture finer variations, each shape is further categorized into three classes based on their dynamic ranges, resulting in a total of 31 pitch contour classes. The dynamic ranges of the pitch frequency are represented with an S (small), M (medium), or B (big) prefix, corresponding to a range of 10–60 Hz, 60–100 Hz, and above 100 Hz, respectively. If the dynamic range is less than 10 Hz, the contour is classified as flat. The basic pitch contour shapes (except flat) are portrayed in Fig. 3.
@@
III Performance Evaluation
III-A Speech Corpus
The corpus used in the current work consists of read speech data recorded by professional, native speakers of Tamil and Hindi. The data for each language comprises around 4000 utterances, each of length up to 3 seconds, recorded by one male and one female speaker, in a studio environment, at a sampling rate of 16 kHz. This adds up to around 12 hours of data in each language for each speaker [17]. Indian English data is recorded from the two native Tamil speakers, also spanning 12 hours each. Around five minutes of data from this corpus for each language is manually segmented at the phoneme level by open source software. Pitch contour labels at the syllable level and break indices are also manually assigned for this data. These are used to evaluate the performance of the proposed algorithms. The remaining data are used to train the phoneme HMMs to perform segmentation.
III-B Accuracy of Segmentation

### v2

@@ Abstract
The availability of prosodic information from speech signals is useful in a wide range of applications. However, deriving this information from speech signals can be a laborious task involving manual intervention. Therefore, the current work focuses on developing a tool that can provide prosodic annotations corresponding to a given speech signal, particularly for Indian languages. The proposed AutoProsody tool provides time-aligned phoneme, syllable, and word transcriptions, syllable-level pitch contour annotations, break indices, and syllable-level relative intensity indices. The tool focuses more on syllable-level annotations since Indian languages are syllable-timed [1]. Indians, regardless of the language they speak, may exhibit influences from other languages. As a result, other languages spoken in India may also exhibit syllable-timed characteristics. The accuracy of the annotations derived from the tool is analyzed by comparing them against manual annotations and the tool is observed to perform well. Although the current work focuses on five languages, namely Tamil, Hindi, and Indian English, Malayalam and Gujarati the tool can be extended to other Indian languages and possibly other languages in syllable time as well.
keywords:
Prosody, Speech Annotation, Indian Languages
1 Introduction
Prosody is a broad term encompassing multiple features such as tones, intonation patterns, pitch accents, and rhythm. In this work, we use the term prosody specifically to refer to measurable acoustic-prosodic correlates: Pitch (F0 contour), Intensity (energy), Break Indices (pause-duration cues). While recent advancements in artificial intelligence have resulted in high-performing speech-based systems, these still do not accommodate prosodic variations in spontaneous speech or discourse. For instance, current text-to-speech (TTS) synthesizers produce speech that is highly intelligible and natural. However, these require a huge amount of training data and are still devoid of emotion and adequate prosodic variations. Incorporating prosodic information while training a TTS system could possibly improve the naturalness of synthetic speech, without necessitating an increase in the amount of training data. For instance, without prosodic cues, the sentence “I can’t believe it” sounds flat and monotone; with prosody, it exhibits natural pitch movement and stress that convey genuine surprise.
Prosodic features influence speech processing in distinct ways:
  • Pitch (F0) distinguishes interrogatives from declaratives and conveys speaker emotion.
  • Relative intensity (RI) correlates with stress and prominence, also supporting emotion recognition.
  • Break Indices (BI) classify inter-word silences into short, medium, and long pauses reflecting boundary strength.
Example 1 — Pitch: “You’re coming.” vs. “You’re coming?” Although the words are identical, a rising pitch contour in the second utterance signals a question.
Example 2 — Intensity: “I didn’t say he stole the money.” vs. “I didn’t say he stole the money.” Shifting intensity (stress) changes the emphasized word and alters the implied meaning.
Example 3 — Break Indices: “Old men $|$ and women were seated.” (medium BI between men and women) vs. “Old men and women were seated.” (short BI between men and women). The pause changes how the sentence is spoken and understood.
@@ 1 Introduction
Another application that can benefit from the availability of prosodic information is speech-to-speech (S2S) translation. An S2S system involves an automatic speech recognition (ASR) system that provides text corresponding to a given speech signal. This text is then translated to the desired language, which is then synthesized using a TTS system. Incorporation of prosodic information along with the ASR output could aid in producing speech in the target language that carries the same tones information as the speech in the source language. This is precisely where we have a problem. Also when the order or sequence of word changes, incorporation of prosody in the target language is a big challenge. Further, each language has a unique prosodic structure and hence prosodic features could aid in language identification and in the incorporation of language specific prosody in TTS and S2S systems as well.
@@
While prosody is immensely useful in a wide range of applications, obtaining accurate prosodic annotations is usually a laborious task requiring manual intervention. Further, while a standard representation of prosody exists for English, namely ToBI (Tones and Break Indices) [2, 3], as well as for certain other languages, no such standard has been established for Indian languages. Therefore, the current work extends on the existing ToBI standard while incorporating additional information, namely time-aligned phonetic, syllabic, and word transcriptions and a relative intensity index, to develop a tool for prosodic annotation.
@@
The ToBI system is a set of conventions for transcribing and annotating the prosody of speech, originally designed for English. Since the manual annotation of prosody is laborious, certain automatic labeling tools have been developed for English. One such tool is the Automatic ToBI (AuToBI) [4], which uses machine learning to predict ToBI labels from acoustic features of the speech signal. Another commonly used platform for annotation is praat [24]. Although originally designed for manual annotation, it hosts several plugins and scripts supporting semi-automatic prosodic annotation [5]. Web MAUS [6] is a web-based service for automatic segmentation and labeling of speech, including prosodic features. Another popular tool is PyToBI [7], which provides a Python interface for automatic labeling with ToBI. However, these are specifically designed for English.
@@
Efforts to study prosody in Indian languages—characterized by their syllable-centric structures—have largely focused on analyzing intonation, stress, and rhythm in languages such as Hindi, Bengali, Tamil, Kannada, and Telugu [15, 16, 17]. However, a standardized framework for representing prosodic features across these languages is yet to be established. Recognizing the similarities in their prosodic structures, this work proposes a common prosodic annotation standard and tool. The tool automates the annotation of pitch contours labels, and break indices for Indian languages, while incorporating additional features, including:
@@
We propose an automatic prosody tool specifically designed to extract prosodic features from Indian language speech. Unlike existing ToBI-based systems, this tool does not aim to assign categorical ToBI labels. Instead, it focuses on extracting continuous prosodic features—such as pitch contours labels, energy, duration measures, and syllable-level timing—which are more flexible and suitable for analyzing the diverse prosodic structures found in Indian languages. The tool is modular, language-agnostic, and scalable, enabling both exploratory prosodic analysis and downstream applications in speech synthesis, recognition, and corpus annotation.
The paper is organized as follows: Section 2 provides a detailed description of the proposed tool, Section 3 evaluates the accuracy of the annotations provided by the tool, Section 4 presents the use of the tool to perform language identification, and Section 5 concludes the paper.
2 The AutoProsody Tool
The AutoProsody tool is designed to provide rich prosodic information from a speech signal. Given a speech signal, the tool converts it to a single channel and resamples it to a sampling frequency of 16 kHz. It then uses an ASR to obtain the corresponding orthographic transcription. This transcription, along with the speech signal, is used to derive phoneme, syllable and word-level boundaries. The tool also processes the speech signal to estimate the relative intensity index, pitch contour labels, and break indices. The extracted information is then plotted as shown in figure 1.
Figure 1: Annotations from the AutoProsody Tool for the English sentence, ”For the twentieth time that evening the two men shook hands.”
In the AutoProsody tool, an automatic speech recognition (ASR) system is used when only the speech signal (WAV format) is available, without accompanying text. In this scenario, the ASR generates an orthographic transcription from the audio. The other parameters are extracted as mentioned above. However, it’s important to note that the transcription produced by the ASR may sometimes be inaccurate, potentially leading to errors in the segmentation of the speech signal and in the extraction of other features.
@@ 2 The AutoProsody Tool
Importantly, in both cases—whether the ASR system is used (when only audio is available) or not (when both audio and text are provided)—the AutoProsody tool consistently applies language-independent models.
Considering the number of languages in India, building systems for each language individually is a complex and time-consuming task. For instance, creating a Text-to-Speech (TTS) system for a new language requires a thorough understanding of its phonotactics (sound structure), letter-to-sound (LTS) rules, and other linguistic elements. This process involves considerable effort and often requires the expertise of linguists familiar with the language. As Prakash et al. (2014) highlighted in their work, ”An approach to building language-independent text-to-speech synthesis for Indian languages” [22],[25], building such systems from scratch for each language involves challenges such as training Hidden Markov Models, developing speech context modeling, and collecting language-specific data. These tasks are not only complex but also time-consuming.
By adopting language-independent models, the AutoProsody tool eliminates the need for such language-specific efforts. These models offer the flexibility of adapting to new languages with minimal modification, making the process more efficient. They do not require the extensive linguistic resources that traditional, language-dependent models need. Instead, they rely on generalizable algorithms that can be easily applied across multiple languages, greatly improving the scalability and applicability of the AutoProsody tool for the multilingual context of India.
@@
The block diagram of AutoProsody tool figure 2 outlines the overall process of text and speech processing, where input speech and text are segmented and are processed via prosody modules.
Figure 2: Block diagram of AutoProsody Tool
2.1 Automatic Speech Recognition
To obtain the orthographic transcription corresponding to a given speech signal, the speech signal and the language (obtained from the user) are given to an ASR. In the current work, a pre-trained Data2Vec-aqc-based ASR model [18], trained on 30,000 hours of speech data from 24 Indian languages, including Tamil, Hindi, and Indian English, Malayalam and Gujarati is employed.
2.2 Segmentation
Once the text corresponding to a given speech signal is obtained from the ASR, it is used to generate time-aligned phonetic, syllabic, and word transcriptions. To achieve this, context-independent hidden Markov models (HMMs) are trained for each phoneme across the five languages considered  [^1] Due to the requirement for less data and faster training, we favor HMM-based training over neural network methods.. Phoneme segmentation is performed using HMM-based forced alignment. The input speech is first converted into acoustic features, and the corresponding phoneme sequence is obtained using a grapheme-to-phoneme (G2P) converter. Given the phoneme sequence and trained HMM acoustic models, forced alignment is applied to determine the start and end time of each phoneme in the speech signal. These HMMs are trained using three states, with one to five mixture components per state, depending on the number of available examples for each phoneme. The forced-Viterbi alignment procedure [19, 23] is initially applied to identify phoneme boundaries, and this process is repeated over four iterations to obtain the final boundaries. This iterative process ensures the accurate segmentation of speech data, enabling robust phoneme, syllable, and word-level boundary derivation. Monophone HMMs for all phonemes are generated using the label files obtained. With these models, forced-Viterbi alignment is performed iteratively to segment the rest of the data, as described below:
@@ 2.2 Segmentation
This iterative process ensures the accurate segmentation of speech data, enabling robust phoneme, syllable, and word-level boundary derivation. Since the tool is designed to support all five languages, language-independent phoneme models are also trained, where acoustically similar phonemes across the languages share the same model.
Syllabification involves combining phoneme boundaries to define syllable boundaries using a set of predefined rules. In this process, each vowel acts as the nucleus of a syllable, with consonants appearing at the beginning or end. The phoneme sequence is grouped into syllables in forms such as $V$, $C^{*}V$, $VC^{*}$, and $C^{*}VC^{*}$. This systematic approach ensures precise syllable segmentation, facilitating subsequent linguistic and prosodic analysis.
While deriving the phoneme boundaries, it is ensured that there are silences inserted at the end of each word, though their duration might be close to zero at several places. These silences are used to combine the phoneme/syllable boundaries to derive word boundaries as discussed in Section 2.2.
2.3 Computation of the Relative Intensity Index
The tool then calculates the relative intensity index at the syllable level. The speech signal is divided into overlapping frames of 20 ms with a 10 ms hop length, and the Short-term energy (STE) is computed for each frame. The energy is subsequently normalized across the entire utterance. Based on the empirical analysis for this normalized energy, a relative intensity index with difference of 0.2 is assigned to each syllable on a scale from 1 to 5 as follows:
@@ 2.3 Computation of the Relative Intensity Index
Fig. 1 (b) depicts $RII$, where the syllable, “mehn” has the highest RII of 5, while the syllables, “dha”, “iyv”, “ningh”, “shuck”, “haendz”, etc. have the lowest RII of 1.
2.4 Estimation of Break Indices
In order to determine the break indices, speech versus silence discrimination is first performed using spectral flatness, which is a measure of the energy distribution across frequencies. Spectral Flatness (SF) is computed frame-wise using the following equation, with the same frame length and hop length mentioned in the previous section.
@@ 2.4 Estimation of Break Indices
[…]
Here, Break Index values 1, 2, and 3 correspond to short, medium, and long inter-word silences, respectively.
2.5 Labeling Pitch Contours
@@
Once the pitch contour, i.e., the time-varying $F_{0}$, is estimated, it undergoes syllable-level smoothing to reduce local fluctuations and abrupt changes. This is achieved using a polynomial of order 3, which fits a smooth curve to the pitch values over time, yielding a more natural representation of prosodic variations. Initially, eleven shapes of the pitch contour, namely L, H, HLL, HHL, LLH, LHH, HLH, LHL, Hat,Bucket, and flat, are identified, where L (low) represents a falling contour and H (high) represents a rising contour. To capture finer variations, each shape is further categorized into three classes based on their dynamic ranges, resulting in a total of 31 pitch contour classes. The dynamic ranges of the pitch frequency are represented with an S (small), M (medium), or B (big) prefix, corresponding to a range of 10–60 Hz, 60–100 Hz, and above 100 Hz, respectively. If the dynamic range is less than 10 Hz, the contour is classified as flat. The basic pitch contour shapes (except flat) are portrayed in figure 3.
@@
3 Performance Evaluation
3.1 Speech Corpus
The speech corpus used for analysis consists of read speech data from both text-to-speech (TTS) and automatic speech recognition (ASR) resources for five Indian languages: Indian English, Tamil, Hindi, Gujarati, and Malayalam. For each language, a total of 6 hours of data is used, comprising 4 hours for training and 2 hours for evaluation.
The training data (4 hours per language) is constructed as follows: (i) 1 hour of clean speech obtained from the IITM IndicTTS corpus [8], (ii) 1 hour of clean speech drawn from the IITM INX-SpeakerHub (multispeaker ASR) corpus [9], and (iii) 2 hours of noise-augmented speech, created by adding car, babble, street, and white noise (0.5 hours each) at 0, 5, and 10 dB SNR levels to the IndicTTS corpus. These noises are from the NOISEX-92[27] and the NOIZEUS datasets[28]. The noise augmentation is applied to assess the robustness of the proposed system under non-ideal acoustic conditions.
The IndicTTS data is recorded by professional male and female speakers in a controlled studio environment to maintain consistent voice quality and pitch characteristics. The ASR data originates from the INX-SpeakerHub corpus, which contains speech from a large number of native Indian speakers across multiple languages and provides substantial speaker diversity. All recordings are mono-channel, sampled at 48 kHz, with 16-bit resolution, and stored in WAV format. This corpus design enables systematic evaluation across multiple languages, speaker conditions, and acoustic environments, while remaining suitable for low-resource prosody annotation tasks.
3.2 Accuracy of Segmentation